# Break-for-Make: Modular Low-Rank Adaptations for Composable Content-Style Customization

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; isbn: ;
@@ 1. Introduction
In this work, we introduce a two-stage learning approach for customized content-style generation, namely “break-for-make”. In the first stage, we propose “partly learnable projection” (PLP) matrices to train content and style in separated sub-parameter spaces of low-rank adapters. Specifically, we freeze certain parameters in both the “up projection” and “down projection” matrices, allowing separate training of content and style within their respective trainable parameter subsets. To avoid interference between content and style after matrix multiplication by frozen parameters, we initialize the frozen rows and columns within the projection matrices to approximate orthogonal bases. To maintain the generalization of the learned content/style PLPs, we utilize a multi-correspondence projection (MCP) learning strategy to learn unbiased content and style parameter spaces. Specifically, we train customized content in “up projection” matrices with diverse reference styles in “down projection” matrices and vice versa. This approach avoids one-to-one binding between content and style, thereby mitigating the overfitting of content/style PLPs when composing with other corresponding PLPs. In the second stage, we reconstruct the unified parameter space using the content and style PLP matrices trained in the first stage, then fine-tune the combined adapter to obtain content-style customized results. As the specific content and style are learned separately and in a generalized manner during the first stage, fine-tuning (approximately a few dozen steps) is required for the combined adapter to generate images that better align with the content and style references, as shown in Figs. 1(a) and (b).
@@ 2.2. Customized Content Style Fusion
The goal of content-style customization is to generate images that effectively present both user-specified content and style while ensuring unique characteristics of both are distinctively represented (Zhang et al., 2023b; Shah et al., 2023). Formally, given a content reference image $I_{c}$, a style reference image $I_{s}$, and a prompt $P$, we aim to generate an output image $I_{out}$ that contains the same content as $I_{c}$, and has the same style as $I_{s}$, while aligning with the provided prompt $P$.
@@ 5.1. Main Results
We compare our method with five content-style customization methods: DreamBooth+LoRA (DB+LoR-A), Textual Inversion (TI), ProSpect, Custom Diffusion (CD) and ZipLoRA. We first present the results of generating the same content image with multiple style images in Fig. 5, then we present the same style image with multiple content images in Fig. 6.
@@
Results indicate that our methods successfully disentangle content and style in one image while maintaining a high level of fidelity. The DB+LoRA method usually generates images of unnatural content style fusion (the result of “mountain” with “yarn style” and “oil painting style” in Fig. 5) and images of the mixed style (“vase” and “teapot” with “glass style” in Fig. 6), the observed entanglement phenomenon aligns with the analysis presented in Section 3. The TI method only updates parameters in the text embedding space, thus having a relatively weaker learning capability. At times, it struggles to accurately learn content/style features, leading to a decrease in the fidelity of generated images (“mountain” with “Minimalism painting style” and “marble style” in Fig. 5, the loss of feature “transparent glass” in “glass style” in Fig. 6). ProSpect learns the reference image with a specific token and trains the embedding of this token, then adds it as a condition in different steps during inference. This method has achieved effective control over content and style to some extent, as seen in examples such as “vase” and “teapot” in Fig. 6, shape and material are presented in generated images. However, it is constrained by its learning capability, which leads to low-quality content-style customization results (the result of “mountain” with “watercolor painting style” and “yarn style” in Fig. 5). The CD also encounters entangling issues between content and style. In cases of “glass style” with “vase” and “teapot” in Fig. 6, the reference images of content influence the style of the generated images. In the case of ZipLoRA, the generated results may not accurately present the reference content or style. For example, in instances like “vase” and “teapot” in Fig. 6, the outputs of ZipLoRA lack the texture style of “transparent glass” in the reference set. In instances of generating “mountain” with “oil painting” style and “blackborad painting” style, the mountain cannot be generated faithfully as the reference. This also reflects the manifestation of fidelity degradation due to parameter conflicts. Compared with the above methods, our method maintains a high level of fidelity and harmonious content-style interaction when generating various styles for the same content. This also demonstrates the strong generalization capability of our approach. One more interesting thing is that the instance of the “sticker style” images includes a dual style, encompassing both sticker and cartoon styles. When evaluating it as the reference, our method successfully generates images in the sticker style. It simultaneously transfers the content into a cartoon style, while the results of other methods are kept in a realistic style.
@@
We present quantitative comparison results in Table 1. We present the style-alignment and prompt-alignment metrics calculated using CLIP feature extraction, as well as the content-alignment metric obtained through DINO feature extraction. Additionally, we showcased the average metric for these three, where higher values indicate better performance. Overall, our average score is the highest among all the baselines, indicating that our method generates customized content-style images that align with the content and style reference and correspond to the prompt. Note that in the content-alignment metric, our score is not the highest because other methods tend to generate images that retain more features from the content reference images. However, this may impact the accurate expression of style and prompt in the generated images, affecting the effectiveness of style generation and prompt alignment. The lower style-alignment and prompt-alignment metrics for other methods in Table 1 also indicate this. Additionally, the comparative display in Fig.5 and Fig. 6 supports this observation.
@@ 5.2. Editability Evaluation and Comparison
We evaluate and compare the editability of our method with other baselines by utilizing a diverse set of prompts to generate customized content-style fusion images. For a fair comparison, the prompts and results of ZipLoRA are from their original paper. As illustrated in Fig. 8, in most cases, ZipLoRA is able to generate customized content-style images effectively and aligns well with the provided prompts. However, in some details, ZipLoRA tends to lose the characteristics of the reference image, such as the ears and mouth in “wearing a hat”, and the overall appearance in “in a boat” and “driving a car”. In contrast, our method maintains consistency with the reference image in these generated results. We additionally showcased more generation results from diverse prompts in the bottom two rows of Fig. 8. Results are highly aligned with the prompts, maintaining a high level of disentanglement between content and style, as well as content and style fidelity.
@@ 5.4. Concept and Style Reconstruction after First Stage Training
In the two-stage training paradigm we propose, the first stage employs a multi-correspondence projection learning strategy that accurately learns the specific content and style features. Consequently, in the second stage, only a few dozen iterations of fine-tuning are required to generate customized content-style fusion images. In Fig. 10, we illustrate that our approach has accurately learned features of content or style in the first stage of training and maintained a high level of fidelity of individuals after the second stage of fine-tuning. Moreover, this multi-correspondence projection learning strategy prevents overfitting between content and style, thereby enabling the generation of more diverse results based on prompts. We also present images generated from directly combined adapters without fine-tuning in Fig. 10; results verify that content and style are disentangled in the first stage and have better effects after fine-tuning.
@@ 5.5. Comparison with Other Two-Stage Content-Style Customization Paradigms
For the task of customized content-style image generation, we also evaluate other two-stage approaches: learning specific content/style in the first stage and subsequently learning or editing style/content (Bau et al., 2019; Hertz et al., 2022; Avrahami et al., 2023b, c; Brooks et al., 2023; Mokady et al., 2023; Kawar et al., 2023; Parmar et al., 2023; Balaji et al., 2022) based on the previous results in the second stage. In our experiments, we learn the content of reference images in the first stage and learn or edit style in the second stage. We leverage NULL-text Inversion  (Mokady et al., 2023), a SOTA real image editing method to edit style in the second stage. Results are presented in Fig. 11. We observe that in the results of the two-stage training/editing methods, both of these methods share similar drawbacks. Specifically, there is entanglement between content and style features. For instance, when generating the “glass” style, “teddy bear” retains the plush features, and “vase” and “teapot” retain opaque material from content reference. In the instance of the “sticker” style, these two methods only generate the contours as the “sticker” style, while the content of the sticker still reflects the realistic style depicted in the content reference image. Besides, the editing-based approach often necessitates complex prompts to accurately describe the features of the reference image, thereby increasing the difficulty of precisely customizing content-style generation. Compared with both methods, our method effectively disentangles the content and style of the reference image, blending them together to generate high-quality customized content-style images without the need for complex prompts.
@@ 5.7. Ablation Study
Figure 13. Analysis of three main methods using LoRA for customized content-style image fusion. Joint training LoRA will mix the parameter space of content and style, leads to entanglement of content and style. Merging LoRAs after independent training has problem of conflict parameters from content LoRA and style LoRA, leads to content and/or style unfaithful after fusion. Our method trains content and style in separated parameter subspaces of LoRA with orthogonal parameter in fixed parts of parameter space, result in entangled and faithful content and style fusion.
In Section 4, we introduce the hyperparameter $d$ as the row dimension of the fixed parameters, representing the proportion of fixed parameters in the parameter subspace. We conduct experiments on eight different fixed parameter ratios, $1/8$, $1/4$, $3/8$, $1/2$, $5/8$, $3/4$, $7/8$, and $1$, representing the proportion of fixed parameters in relation to the total parameters. We quantitatively test the text alignment, content alignment, and style alignment metrics for various $d$, and the results are presented in Fig. 13. From the histogram, we observe that as the ratio of fixed parameters increases, both values of “Style Alignment” and “Text Alignment” gradually rise, reaching their peaks at a ratio of $0.5$, and then gradually decline. The “Average Alignment” reaches its peak at a ratio of $0.5$. This indicates that the optimal alignment occurs at a ratio of $0.5$, resulting in better customized content-style images. This finding aligns with our theoretical framework introduced in Section 4, where a 1:1 ratio between fixed and trainable parameters results in the “content parameter subspace” in Eq. (12) and “style parameter subspace” in Eq. 11) having the maximum number of trainable parameters, thus reaching the maximum learning capacity and achieving the best generation effect. Note that at a ratio of $0.5$, the “Content Alignment” is not maximal. This is because the results of other ratios present a weaker learned style (as indicated by lower “Style Alignment” in the histogram) and are entangled with the content to some extent.
Figure 14. Ablation study evaluating the impact of the proposed Multi-Correspondence Projection (“MCP”). We train specific content (e.g.“vase”) and style (e.g.“glass”) in a one-to-one manner and directly inference after training. Results are presented in W/o MCP - I column. We train specific content (e.g.“vase”) and style (e.g.“yarn”) in a one-to-one manner in the first stage, and combine the content (e.g.“vase”) adapters with other style (e.g.“glass”) adapters in the second stage, then inference with the combined adapters. Results are presented in W/o MCP - II column. The visual comparison highlights the effectiveness of MCP in enhancing the details while preserving the disentanglement of content and style, as well as maintaining high-level fidelity of them.
@@
To prevent overfitting between specific content and style during the training stage, we introduce multi-correspondence projection (“MCP”) learning within our work. We conduct two ablation studies with different experiment settings for evaluating the impact of the proposed MCP. Specifically, in the first study, we train the specific content (e.g., “vase”) with style (e.g., “glass style”) in a one-to-one manner and leverage the trained model for inference; results are shown in W/o MCP-I in Fig. 14. In the second study, we train the specific content (e.g., “vase”) with another style (e.g., “yarn style”) in a one-to-one manner in the first stage and combine the content adapters with the trained style (e.g., “glass style”) adapters in the second stage, at last, we use the combined model for inference, results are shown in W/o MCP-II in Fig. 14. We can observe that in the first study, the generated images exhibit a degree of overfitting to the reference images(“teapot” with “chair legs” from the style reference image, “sticker” style with realistic style from the content reference image), resulting in a decrease in fidelity to the content or style, thereby reducing the quality of the outputs. In the second study, we can observe that the results exhibit some features (the color from “yarn style”) of the style trained in the first stage. As the final model does not incorporate this style, this is mainly because without MCP, the “yarn style” influences the parameter space of the content during the training stage, as analyzed in Fig. 4. We present the results of our methods in With MCP; by comparing, we can observe that with MCP, we can effectively avoid overfitting and generate images with more disentangled content and style.
@@
To demonstrate the effectiveness of the orthogonal fixed parameters designed to enhance the content and style fidelity of generated images, we conducted an experiment where we removed the orthogonal fixed parameters and replaced them with randomly fixed parameters. We present results in Fig.15 for comparison. Without the orthogonality of the fixed parameters, it leads to decreased fidelity for generated images. For instance, in “vase”, “teapot”, and “teddy bear”, the generated images no longer preserve the original content details, and the style also changed. In the case of the “sticker style”, the generated images lose the cartoonish style of contents present in the reference. We also present quantitative results in Table. 1, after ablating fixed parameter orthogonalization, although the content alignment slightly increases, the style alignment decreases significantly, and the average alignment decreases. Note that the slight increase in text alignment is due to the increase in content alignment.
@@
In the second stage of our pipeline, we reconstruct the entity parameter space by the content and style PLP matrices. Then, we fine-tune the combined LoRA modules for a few dozen steps to make the model generate images of customized content with style. We present the results of the ablating fine-tuning step in Fig. 10. The results demonstrate that after undergoing a few dozen fine-tuning steps, our proposed method achieved optimal visual performance. We also present individual content or style generation results in the middle and bottom rows of Fig. 10. These results demonstrate that our proposed method successfully disentangles content and style while retaining the capability to faithfully generate individual content or style.
@@
We want to evaluate the learning effect of the desired content or style compared with the basic stable diffusion model. We use pseudo words for training and inference specific content and style. For comparison, we describe content and style in prompt for generation. As results present in Fig. 16, only leverage prompt to describe the desired content or style and without learning them cannot capture detailed features from reference images, leads to unfaithful generation of the customized content and style.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; ;
@@ 1. Introduction
In this work, we introduce a two-stage learning approach for customized content-style generation, namely “break-for-make”. In the first stage, we propose “partly learnable projection” (PLP) matrices to train content and style in separated sub-parameter spaces of low-rank adapters. Specifically, we freeze certain parameters in both the “up projection” and “down projection” matrices, allowing separate training of content and style within their respective trainable parameter subsets. To avoid interference between content and style after matrix multiplication by frozen parameters, we initialize the frozen rows and columns within the projection matrices to approximate orthogonal bases. To maintain the generalization of the learned content/style PLPs, we utilize a “multi-correspondence projection” (MCP) learning strategy to learn unbiased content and style parameter spaces. Specifically, we train customized content in “up projection” matrices with diverse reference styles in “down projection” matrices and vice versa. This approach avoids one-to-one binding between content and style, thereby mitigating the overfitting of content/style PLPs when composing with other corresponding PLPs. In the second stage, we reconstruct the unified parameter space using the content and style PLP matrices trained in the first stage, then fine-tune the combined adapter to obtain content-style customized results. As the specific content and style are learned separately and in a generalized manner during the first stage, fine-tuning (approximately a few dozen steps) is required for the combined adapter to generate images that better align with the content and style references, as shown in Fig. 1(a) and (b).
@@ 2.2. Customized Content Style Fusion
The goal of content-style customization is to generate images that effectively present both user-specified content and style while ensuring their unique characteristics are distinctively represented (Zhang et al., 2023b; Shah et al., 2023). Formally, given a content reference image $I_{c}$, a style reference image $I_{s}$, and a prompt $P$, we aim to generate an output image $I_{out}$ that contains the same content as $I_{c}$, and has the same style as $I_{s}$, while aligning with the provided prompt $P$.
@@ 5.1. Main Results
We compare our method with five content-style customization methods: DreamBooth+LoRA (DB+LoR-A), Textual Inversion (TI), ProSpect, Custom Diffusion (CD) and ZipLoRA. We first present the results of generating the same content image with multiple style images in Fig. 5, then we present the same style image with multiple content images in Fig. 6. Results indicate that our methods successfully disentangle content and style in one image while maintaining a high level of fidelity. The DB+LoRA method usually generates images of unnatural content style fusion (the result of “mountain” with “yarn style” and “oil painting style” in Fig. 5) and images of the mixed style (“vase” and “teapot” with “glass style” in Fig. 6), the observed entanglement phenomenon aligns with the analysis presented in Section 3. The TI method only updates parameters in the text embedding space, thus having a relatively weaker learning capability. At times, it struggles to accurately learn content/style features, leading to a decrease in the fidelity of generated images (“mountain” with “Minimalism painting style” and “marble style” in Fig. 5, the loss of feature “transparent glass” in “glass style” in Fig. 6). ProSpect learns the reference image with a specific token and trains the embedding of this token, then adds it as a condition in different steps during inference. This method has achieved effective control over content and style to some extent, as seen in examples such as “vase” and “teapot” in Fig. 6, shape and material are presented in generated images. However, it is constrained by its learning capability, which leads to low-quality content-style customization results (the result of “mountain” with “watercolor painting style” and “yarn style” in Fig. 5). The CD also encounters entangling issues between content and style. In cases of “glass style” with “vase” and “teapot” in Fig. 6, the reference images of content influence the style of the generated images. In the case of ZipLoRA, the generated results may not accurately present the reference content or style. For example, in instances like “vase” and “teapot” in Fig. 6, the outputs of ZipLoRA lack the texture style of “transparent glass” in the reference set. In instances of generating “mountain” with “oil painting” style and “blackborad painting” style, the mountain cannot be generated faithfully as the reference. This also reflects the manifestation of fidelity degradation due to parameter conflicts. Compared with the above methods, our method maintains a high level of fidelity and harmonious content-style interaction when generating various styles for the same content. This also demonstrates the strong generalization capability of our approach. One more interesting thing is that the instance of the “sticker style” images includes a dual style, encompassing both sticker and cartoon styles. When evaluating it as the reference, our method successfully generates images in the sticker style. It simultaneously transfers the content into a cartoon style, while the results of other methods are kept in a realistic style.
@@
We present quantitative comparison results in Table 1, evaluating the style-alignment, prompt-alignment (using CLIP feature extraction), and content-alignment (using DINO feature extraction) metrics. Additionally, we report the average of these three metrics, where higher values indicate better performance. Our method achieves the highest average score among all baselines, suggesting it generates customized content-style images that align well with the content and style references while corresponding to the given prompt. Note that in the content-alignment metric, our score is not the highest because other methods tend to generate images that retain more features from the content reference images. However, this could compromise accurate expression of style and adherence to the prompt in the generated images, affecting the effectiveness of style transfer and prompt alignment, as indicated by the lower style-alignment and prompt-alignment metrics for other methods in Table 1. Additionally, the comparative display in Fig.5 and Fig. 6 supports this observation.
@@ 5.2. Editability Evaluation and Comparison
We evaluate and compare the editability of our method against other baselines by generating customized content-style fusion images using a diverse set of prompts. For a fair comparison, the prompts and results of the ZipLoRA are obtained from their original paper. As illustrated in Fig. 8, ZipLoRA is generally effective in generating customized content-style images that align well with the provided prompts. However, in some details, ZipLoRA tends to lose certain characteristics of the reference image, such as the ears and mouth in the “wearing a hat” example, and the overall appearance in the “in a boat” and “driving a car” examples. In contrast, our method maintains better consistency with the reference image in these generated results. Additionally, we showcased more generation results from diverse prompts in the bottom two rows of Fig. 8. These results demonstrate high alignment with the prompts while maintaining a high level of disentanglement between content and style, as well as preserving the fidelity of content and style representations. Overall, our method exhibits superior editability compared to existing baselines, enabling the generation of customized content-style images that faithfully integrate the provided prompts while retaining the desired characteristics of the reference content and style.
@@ 5.4. Concept and Style Reconstruction after First Stage Training
The proposed two-stage training paradigm employs a multi-corresp-ondence projection learning strategy in the first stage to accurately learn the specific content and style features. Consequently, in the second stage, only a few dozen iterations of fine-tuning are required to generate customized content-style fusion images. Fig. 10 illustrates that our approach has accurately learned the features of content or style in the first stage of training and maintained a high level of fidelity for individual content and style after the second stage of fine-tuning. Moreover, this multi-correspondence projection learning strategy prevents overfitting between content and style, thereby enabling the generation of more diverse results based on prompts. We also present images generated from directly combined adapters without fine-tuning in Fig. 10, these results verify that content and style are disentangled in the first stage and have better effects after undergoing the fine-tuning process.
@@ 5.5. Comparison with Other Two-Stage Content-Style Customization Paradigms
For the task of customized content-style image generation, we also evaluate other two-stage approaches that involve learning specific content/style in the first stage and subsequently learning or editing style/content (Bau et al., 2019; Hertz et al., 2022; Avrahami et al., 2023b, c; Brooks et al., 2023; Mokady et al., 2023; Kawar et al., 2023; Parmar et al., 2023; Balaji et al., 2022) based on the previous results in the second stage. In our experiments, we learn the content of reference images in the first stage and learn or edit style in the second stage. We leverage NULL-text Inversion  (Mokady et al., 2023), a SOTA real image editing method to edit style in the second stage. The results are presented in Fig. 11. We observe that both the two-stage training and editing methods share similar drawbacks, primarily the entanglement between content and style features. For instance, when generating the “glass” style, the “teddy bear” retains plush features, and the “vase” and “teapot” retain opaque material from the content reference. In the case of the “sticker” style, these two methods only generate the contours as the “sticker” style, while the content of the sticker still reflects the realistic style depicted in the content reference image. Furthermore, the editing-based approach often necessitates complex prompts to accurately describe the features of the reference image, thereby increasing the difficulty of precisely customizing content-style generation. In contrast, our method effectively disentangles the content and style of the reference image, blending them together to generate high-quality customized content-style images without the need for complex prompts. Our approach demonstrates superior performance in achieving faithful content-style fusion compared to both the two-stage training and editing methods.
@@ 5.6. User Study
    Figure 13. Cosine similarity between features of output and reference style, content and prompt of different dimension d. When d=0.5, the average cosine similarity of the features reaches its maximum, indicating optimal alignment between the generated results and the reference content, style, and prompt.
@@
Figure 14. Ablation study evaluating the impact of the proposed Multi-Correspondence Projection (“MCP”). We train specific content (e.g., “vase”) and style (e.g., “glass”) in a one-to-one manner and directly inference after training. Results are presented in W/o MCP - I column. We train specific content (e.g., “vase”) and style (e.g., “yarn”) in a one-to-one manner in the first stage, and combine the content (e.g., “vase”) adapters with other style (e.g., “glass”) adapters in the second stage, then inference with the combined adapters. Results are presented in W/o MCP - II column. The visual comparison highlights the effectiveness of MCP in enhancing the details while preserving the disentanglement of content and style, as well as maintaining high-level fidelity of them.
@@ 5.7. Ablation Study
In Section 4, we introduce the hyperparameter $d$ as the row dimension of the fixed parameters, representing the proportion of fixed parameters in the parameter subspace. We conduct experiments on eight different fixed parameter ratios, $1/8$, $1/4$, $3/8$, $1/2$, $5/8$, $3/4$, $7/8$, and $1$, corresponding to the proportion of fixed parameters relative to the total parameters. We quantitatively evaluate the text alignment, content alignment, and style alignment metrics for various values of $d$, and the results are presented in Fig. 13. From the histogram, we observe that as the ratio of fixed parameters increases, both the values of “Style Alignment” and “Text Alignment” gradually rise, reaching their peaks at a ratio of $0.5$, and then gradually decline. The “Average Alignment” reaches its maximum at a ratio of $0.5$. This indicates that the optimal alignment occurs at a ratio of $0.5$, resulting in better customized content-style images. This finding aligns with our theoretical framework introduced in Section 4, where a 1:1 ratio between fixed and trainable parameters results in the “content parameter subspace” in Eq. (12) and “style parameter subspace” in Eq. 11) having the maximum number of trainable parameters, thus reaching the maximum learning capacity and achieving the best generation effect. It is noteworthy that at a ratio of $0.5$, the “Content Alignment” is not maximal. This is because the results of other ratios present a weaker learned style (as indicated by lower “Style Alignment” in the histogram) and are entangled with the content to some extent.
@@
To prevent overfitting between specific content and style during the training stage, we introduce a Multi-Correspondence Projection (“MCP”) learning within our work. We conduct two ablation studies with different experimental settings to evaluate the impact of the proposed MCP. Specifically, in the first study, we train the specific content (e.g., “vase”) with a particular style (e.g., “glass style”) in a one-to-one manner and leverage the trained model for inference. The results are shown as W/o MCP-I in Fig. 14. In the second study, we first train the model on a specific content (e.g., “vase”) with a different style (e.g., “yarn style”) in a one-to-one manner. Subsequently, in the second stage, we combine the content adapters with the trained style (e.g., “glass style”) adapters and utilized the combined model for inference. The results are shown as W/o MCP-II in Fig. 14. We can observe that in the first study, the generated images exhibit a degree of overfitting to the reference images(e.g., “teapot” with “chair legs” from the style reference image, “sticker” style with realistic style from the content reference image), resulting in a decrease in fidelity to the content or style, thereby reducing the quality of the outputs. In the second study, we can observe that the results exhibit some features (e.g., the color from “yarn style”) of the style trained in the first stage. As the final model does not incorporate this style, this is mainly because due to the fact that without MCP, the “yarn style” influences the parameter space of the content during the training stage, as analyzed in Fig. 4. We present the results of our methods in With MCP; by comparing, we can observe that with MCP, we can effectively avoid overfitting and generate images with more disentangled content and style.
@@
To demonstrate the effectiveness of the orthogonal fixed parameters designed to enhance the content and style fidelity of generated images, we conduct an experiment where we remove the orthogonal fixed parameters and replace them with randomly fixed parameters. We present the results in Fig.15 for comparison. Without the orthogonality of the fixed parameters, it leads to decreased fidelity for the generated images. For instance, in the case of “vase”, “teapot”, and “teddy bear”, the generated images no longer preserve the original content details, and the style has also changed. In the case of the “sticker style”, the generated images lose the cartoonish style of the contents present in the reference. We also present quantitative results in Table. 1. After ablating fixed parameter orthogonalization, although the content alignment slightly increases, the style alignment decreases significantly, and the average alignment decreases as well. Note that the slight increase in text alignment is due to the increase in content alignment, as the prompts’ emphasis on describing the image content.
@@
In the second stage of our pipeline, we reconstruct the entity parameter space by combining the content and style PLP matrices. Subsequently, we fine-tune the combined LoRA modules for a few dozen steps to enable the model to generate images with customized content with style. The results of ablating the fine-tuning step are presented in Fig. 10. The results demonstrate that after undergoing a few dozen fine-tuning steps, our proposed method achieves optimal visual performance. We also present individual content or style generation results in the middle and bottom rows of Fig. 10. These results illustrate that our proposed method successfully disentangles content and style while retaining the capability to faithfully generate individual content or style.
@@
We aim to evaluate the learning effect of the desired content or style in comparison with the baseline stable diffusion model. To achieve this, we employ pseudo words for training and inference of specific content and style. For the purpose of comparison, we describe content and style using prompts for generation. As the results presented in Fig. 16, solely relying on prompts to describe the desired content or style, without learning these representations, fails to capture detailed features from reference images, leads to unfaithful generation of the customized content and style.